# Minimizing the number of edges in LC-equivalent graph states

### v1

@@ Appendix D Generalized repeater graph states
To find a better fusion order, we propose applying LC-gates corresponding to the $H$ to $G$ conversion to the resource states before beginning the fusion process. This leads to the following steps: (1) apply LC-gates, (2) follow fusion order to create $H$, and (3) fuse leaves $L$. We observe that fusions commute with each other as they act on different qubits. Therefore, we can combine the construction of $H$ and the addition of leaves $L$ into a single step. This simplifies our protocol into two steps: (1) apply LC-gates to suitable qubits of the resource states, (2) follow the fusion order for creating $H$+$L$ using these modified resource states. The final state from this fusion order is $G$+$L$ due to the LC-gates applied at the beginning.
@@
Commuting the LC-gates to the beginning of the protocol can be understood by stepping away from the graph state picture and considering the circuit of the protocol. Consider the steps in construction of $G$+$L$ following the Intermediate-LC protocol: (1) construct the graph $H$, (2) apply LC-gates to convert $H$ to $G$, and (3) fuse leaves $L$ to the graph $G$. We make the following observations from the circuit picture: (1) the target qubits for LC-gates have not been involved in any fusions until that point in time (otherwise they would have been measured out). As a result, the LC-gates can be pushed to the beginning of the protocol as they commute with the fusions implemented to construct $H$. (2) Pauli corrections may be required on qubits adjacent to those involved in fusions. If an LC-gate acts on such a qubit, the corrections can be modified accordingly. This is possible since the Clifford group normalizes the Pauli group, which allows us to commute the LC-gates through the Pauli corrections. Notably, Ref. [28] also applies LC-gates corresponding to local complementations. These gates must also be accounted for when commuting LC-gates to the beginning.
@@
In Fig. 6, we demonstrate the construction of an RGS by commuting LC-gates to the beginning. We first find the suboptimal fusion order for creating $G$+$L$ using Intermediate-LC, shown in Fig. 6(a). We then construct the fusion network (similar to Ref. [28]) to determine the location of each LC-gate. A fusion network depicts the order in which each resource state has to be fused. We show the fusion network corresponding to Intermediate-LC in Fig. 6(b). The fusion network also details the information about the LC-gates in the context of the required fusions. We use this information to apply the LC-gates to the appropriate qubits of the resource states at the start. This allows us to reorder the fusions to optimize the number of resources.
@@
The LC-gates corresponding to a local complementation operation on a node $v$ correspond to the application of the HSH gate on the qubit corresponding to node $v$ and the S gate to all the qubits that correspond to the neighbors of node $v$. These gates altogether convert a graph state into another graph state. Applying a subset of these gates to the graph state will result in a state that is not a graph state but is LC-equivalent to a graph state: a non-graph state. When we commute LC-gates corresponding to the conversion of graph $H$ to $G$, we distribute these gates among multiple resource states, as shown in Fig. 6(c). Thus, the resource states become non-graph states (but are still LC-equivalent) after the application of these LC-gates. We then fuse these non-graph states following the fusion order for creating $H$+$L$ to obtain our desired graph state of $G$+$L$. Thus, the final equality in Step-4 of Fig. 6(c) takes a non-graph state (LHS) to a graph state (RHS).
@@
In conclusion, the Commute-LC protocol enables the creation of $G$+$L$ by following the procedure for constructing $H$+$L$. Importantly, applying LC-gates to the resource states converts them into LC-equivalent non-graph states. The locations of the LC-gates can be identified from the fusion network for creating $G$+$L$ via Intermediate-LC. This approach has two main advantages: (1) it reduces the number of required fusions and resource states by creating $H$+$L$, and (2) it enables obtaining the optimal fusion order using OptGraphState from Ref. [28].

### v2

@@ Appendix D Generalised repeater graph states
To find a better fusion order, we propose applying LC-gates corresponding to the $H$ to $G$ conversion to the resource states before beginning the fusion process. This leads to the following steps: (1) apply LC-gates, (2) follow fusion order to create $H$, and (3) fuse leaves $L$. We observe that fusions commute with each other as they act on different qubits. Therefore, we can combine the construction of $H$ and the addition of leaves $L$ into a single step. This simplifies our protocol into two steps: (1) apply LC-gates to suitable qubits of the resource states, (2) follow the fusion order for creating HH+$L$ using these modified resource states. The final state from this fusion order is GG+$L$ due to the LC-gates applied at the beginning.
@@
Commuting the LC-gates to the beginning of the protocol can be understood by stepping away from the graph state picture and considering the circuit of the protocol. Consider the steps in construction of GG+$L$ following the Intermediate-LC protocol: (1) construct the graph $H$, (2) apply LC-gates to convert $H$ to $G$, and (3) fuse leaves $L$ to the graph $G$. We make the following observations from the circuit picture: (1) the target qubits for LC-gates have not been involved in any fusions until that point in time (otherwise they would have been measured out). As a result, the LC-gates can be pushed to the beginning of the protocol as they commute with the fusions implemented to construct $H$. (2) Pauli corrections may be required on qubits adjacent to those involved in fusions. If an LC-gate acts on such a qubit, the corrections can be modified accordingly. This is possible since the Clifford group normalises the Pauli group, which allows us to commute the LC-gates through the Pauli corrections. Notably, Ref. [31] also applies LC-gates corresponding to local complementations. These gates must also be accounted for when commuting LC-gates to the beginning.
@@
In Fig. 6, we demonstrate the construction of an RGS by commuting LC-gates to the beginning. We first find the suboptimal fusion order for creating GG+$L$ using Intermediate-LC, shown in Fig. 6(a). We then construct the fusion network (similar to Ref. [31]) to determine the location of each LC-gate. A fusion network depicts the order in which each resource state has to be fused. We show the fusion network corresponding to Intermediate-LC in Fig. 6(b). The fusion network also details the information about the LC-gates in the context of the required fusions. We use this information to apply the LC-gates to the appropriate qubits of the resource states at the start. This allows us to reorder the fusions to optimise the number of resources.
@@
The LC-gates corresponding to a local complementation operation on a node $v$ correspond to the application of the HSH gate on the qubit corresponding to node $v$ and the S gate to all the qubits that correspond to the neighbours of node $v$. These gates altogether convert a graph state into another graph state. Applying a subset of these gates to the graph state will result in a state that is not a graph state but is LC-equivalent to a graph state: a non-graph state. When we commute LC-gates corresponding to the conversion of graph $H$ to $G$, we distribute these gates among multiple resource states, as shown in Fig. 6(c). Thus, the resource states become non-graph states (but are still LC-equivalent) after the application of these LC-gates. We then fuse these non-graph states following the fusion order for creating HH+$L$ to obtain our desired graph state of GG+$L$. Thus, the final equality in Step-4 of Fig. 6(c) takes a non-graph state (LHS) to a graph state (RHS).
@@
In conclusion, the Commute-LC protocol enables the creation of GG+$L$ by following the procedure for constructing HH+$L$. Importantly, applying LC-gates to the resource states converts them into LC-equivalent non-graph states. The locations of the LC-gates can be identified from the fusion network for creating GG+$L$ via Intermediate-LC. This approach has two main advantages: (1) it reduces the number of required fusions and resource states by creating HH+$L$, and (2) it enables obtaining the optimal fusion order using OptGraphState from Ref. [31].
Appendix E Example of MERs
In this section we illustrate the difference between the edge minimisation performance of EDM-SA and EDM-SAILP.
Figure 7: We sample ER graphs $G(n,p)$ with increasing $n$ from 11 to 16 and $p=0.8$. We run EDM-SA and EDM-SAILP on these graphs, with $k_{\textrm{max}}$ and $T(1)$ set to 100. We illustrate a set of three graphs, namely: (1) input graphs, (2) approximate MERs obtained from EDM-SA and (3) the exact MERs obtained from EDM-SAILP. Each row has two such sets with the same number of vertices, and the number of vertices increase from 11 to 16 over all rows. We can observe that both the algorithms provide significant reductions in the number of edges, this is partly be due to the input graph being dense. We also observe that in some cases the number of edges in the approximate MERs obtained from EDM-SA are a little higher than the MERs.